OpenStudy (anonymous):

A copy machine makes 119 copies in 4 minutes and 15 seconds. How many copies does it make per minute?

OpenStudy (anonymous):

First covert the 4 minutes and 15 seconds all to seconds giving you 255 seconds so: \[\frac{119copies}{255\sec}*\frac{60\sec}{\min}=\frac{28copies}{\min}\]